• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 15:59:53 on July21,2003

    Entry number 105893

    This entry is a followup to: 105859

    keyword=Day Shift Summary (II)

    day shift summary continue:

    Beam back at 14:20
    resume data taking at kin5.12, BHWP in, target 270, previous total 7M
    run 2631-2642 are tests to fix left arm DAQ.

    run 2643/22643 14:22-14:36 3M coda events, total 10M
    run 2644/22644 14:41-14:56 3M total 13M
    Stop beam for Sirish to put Compton electron detector in

    run 2645/22645 15:10-15:25 3M total 16M
    run 2646/22646 15:26-15:40 3M total 19M
    15:40 requested a bleedthrough measurement, with Hall A 2.1uA and Hall C 58uA,
    we have 2.29% bleedthrough.
    run 2647/22647 15:42-15:58 3M total 22M